A national survey regarding the "Role of Secretaries in the Information Age" reported that 60% of secretaries eye strain when using computers at work. An executive secretary decided to test this claim as she thought it would be incorrect. She randomly surveyed 50 secretaries and determined that 35 of the experienced eye strain. Based upon her sample data, can she reject the claim at a=.01?

Q. What is the decision rule for this hypothesis test? (You must do the hypothesis test)

a) Reject Ho b) Fail to reject Ho

Write decision rule:

Conclusion:
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is

H0 : p = 0.60

Ha : p 0.60

= x / n = 35 / 50 = 0.70

P0 = 0.60

1 - P0 = 0.40

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

= 0.70 - 0.60 / [(0.60 * 0.40) / 50]

= 1.44

P(z > 1.44) = 1 - P(z < 1.44) = 0.0749

P-value = 0.0749

= 0.01

P-value >

Fail to reject the null hypothesis .

There is not sufficient evidence to suggest that the claim as she thought it would be incorrect .